# Semiparametric Bayesian Difference-in-Differences

### v1

@@ Abstract
Keywords: Difference-in-differences, causal inference, double robustness, nonparametric Bayesian inference, Bernstein–von Mises theorem, Riesz representation, Gaussian processes, Dirichlet processes.
@@ 1 Introduction
We also provide an extension of our Bayesian procedure, which incorporates robustification via estimated propensity scores and is particularly suited for more complex models, either due to a larger number of continuous covariates or when the underlying conditional mean functions are not smooth. Our Double Robust Bayesian procedure adjusts the prior and posterior distributions by incorporating an efficient influence function. By doing so, we leverage the rich frequentist literature on double-robust estimation, specifically the work of Sant’Anna and Zhao (2020) in the DiD framework, without sacrificing many of the desirable properties of the Bayesian approach. Under double-robust smoothness conditions, our robust Bayesian procedure satisfies the semiparametric Bernstein–von Mises (BvM) theorem, albeit with a “bias term” in the posterior. Specifically, the resulting posterior distribution depends on the unknown true conditional mean and propensity score functions. Our double-robust Bayesian approach addresses this “bias term” by incorporating an explicit posterior correction. Both the prior adjustment and the posterior correction are derived from functional forms closely associated with the efficient influence function of estimating the ATT. Due to differences in the efficient influence functions for average treatment effects (ATE) and ATT, our double-robust method shares similarities with but differs in functional form corrections from (Breunig, Liu, and Yu, 2023)’s work on ATE. Notably, we find that the bias term in the BvM for ATT is substantially simpler than that for ATE, which also explains the favorable finite-sample behavior of Bayesian ATT estimators even without posterior corrections.
@@
In our Monte Carlo simulations, we find that our methods result in improved empirical coverage probabilities while maintaining competitive confidence interval lengths compared to existing frequentist methods. This finite sample advantage is also observed in low dimensional cases for our Bayesian method that does not involve prior or posterior corrections. [^1] In contrast, a Bayesian method without prior correction performs poorly for the average treatment effect (ATE) as shown by Breunig, Liu, and Yu (2023). This can be explained by the construction of the Bayesian procedure for the ATT, which involves only a prior specification for the conditional mean function in the control arm. In particular, we note that our approach leads to more accurate uncertainty quantification and is less sensitive to estimated propensity scores that are close to boundary values. Our Bayesian methodology requires a prior specification through a likelihood function for the control arm, for which we impose an exponential family structure. We stress that the misspecification of this structure does not have serious consequences for estimation of the ATT. First, as shown in (Kleijn and van der Vaart, 2006, Section 4), nonparametric Bayesian methods possess the same robustness to misspecification as the frequentist M-estimation using least squares. Second, we provide finite sample evidence through simulations, where we find that our Bayesian procedures are not sensitive to misspecifications of the likelihood functions.
@@
The remainder of this paper is organized as follows. Section 2 presents the setup and introduces the Bayesian framework. Section 3 outlines our Bayesian methods. In Section 4, we establish inference via semiparametric BvM theorems for our first method. In Section 5, we derive a doubly robust, semiparametric BvM theorem for our second method. Section 6 provides BvM results under primitive conditions when using squared expontential process priors. Section 7 presents finite sample results via simulations and an empirical illustration. Towards the end, we outline two extensions of our methodology: Section 8 provides an extension to the staggered intervention with multiple time periods and repeated cross-sectional data. Proofs of main theoretical results are collected in Appendix A. Supplementary Appendices BE provide additional technical results and further simulation evidence.
@@ 2.1 Setup
the conditional mean of the differenced outcome across two periods, where $\Delta Y_{i}:=Y_{i2}-Y_{i1}$. The researcher observes an independent and identically distributed (i.i.d.) observations of $(Y_{i1},Y_{i2},D_{i},X_{i}^{\top})^{\top}$, $i=1,\dots,n$. In addition to this canonical panel data setup, we discuss how our results translate to repeated cross-sections data in Section 8.2, while an extension to staggered intervention is provided in Section 8.1. For notational simplicity, we will henceforth suppress the unit index $i$.
@@
One can construct an estimator that replaces the conditional mean function $m_{0}$ with an estimator, known as the outcome regression approach, as described in Heckman, Ichimura, and Todd (1997). As noted by (Abadie, 2005, p.6), plug-in estimators based on standard nonparametric estimators of the conditional mean function $m_{0}$ can face significant challenges due to the curse of dimensionality. This is where we can capitalize the strength of Bayesian estimation, which allows us to incorporate rich covariate information in the prior distribution. It has also been noted in the recent literature that, in the presence of heterogeneous treatment effects in $X$, i.e., when $\mathbb{E}_{0}[Y_{2}(1)-Y_{2}(0)|X]$ varies with $X$, the two-way fixed effect estimator (TWFE) is not consistent for the ATT. See also Remark 1 of Sant’Anna and Zhao (2020) for an explicit discussion.
@@ 2.2 A Bayesian Framework
where $f_{\Delta Y\mid D,X}(y\mid d,x)$ for $d=1$ is the unrestricted conditional density of $\Delta Y$ given $(D,X)$, while for $d=0$, we impose the exponential family condition as in (2.4). Here, $f$ denotes the joint probability density function of $(D\Delta Y,D,X^{\top})^{\top}$ under $P_{\eta}$, and the corresponding cumulative distribution function is denoted by $F$. Importantly, specifying only the prior distribution on density function $f$ and the conditional mean function $m$ is sufficient for identifying the ATT parameter of interest. Specifically, there is no need to additionally parameterize the propensity score $\pi$, the marginal density of $X$, or the conditional density for the control group, $f_{\Delta Y|D,X}(y\mid 0,x)$, which is a deterministic function of $m(x)$ due to the exponential family assumption imposed in (2.4) below.
We consider the following reparametrization of $(m,f)$ given by $\eta=(\eta^{m},\eta^{f})$. We index the probability model by $P_{\eta}$, where
$m_{\eta}=q^{-1}(\eta^{m})\,\,\text{ and }\,\,f_{\eta}=\exp(\eta^{f}).$
We can write the ATT depending on a hyperparameter $\eta$ as
@@
The family (2.4) allows for counting and continuous outcomes. For instance, when $a=1$, the Poisson distribution corresponds to the choices $c(y)=1/(y!)$, $q(m)=\log m$, and $A(m)=m$, while the exponential distribution is represented by $c(y)=1$, $q(m)=-1/m$, and $A(m)=\log m$. Furthermore, the normal distribution with $\text{Var}(\Delta Y|D=0,X)=\sigma^{2}$ for some $\sigma>0$, is captured by $c(y)=\exp(-y^{2}/(2\sigma^{2}))/\sqrt{2\pi\sigma^{2}}$, $q(m)=m/\sigma$, $A(m)=m^{2}/(2\sigma^{2})$, and $a=1/\sigma$. We note that while generalization to multinomial outcomes, as in Breunig, Liu, and Yu (2023), is possible, we do not consider this case explicitly in this paper.
@@ 3.1 Bayesian Inference using Standard Gaussian Process Priors
We first consider nonparametric Bayesian inference, which builds on a standard Gaussian process prior for the conditional mean function combined with an independent Dirichlet process prior for the conditional expectation in (2.3). The proposed method does not include propensity score adjustment, which prevents it from achieving double robustness, as we see in the next section. However, simulations show that the proposed method is robust to a failure of overlap.
@@
The Dirichlet process is default prior on spaces of probability measures. In accordance with $\mathbb{E}_{\eta}(\cdot)$ in the definition of $\tau_{\eta}$, one can conceptualize that we assign Dirichlet process prior to model the distribution $P_{\eta}$, which induces the Bootstrap weights when the base measure in the Dirichlet process is taken to be zero. This is known as the Bayesian bootstrap (Rubin, 1981) with straightforward implementation; also see Chamberlain and Imbens (2003).
@@
     (a) Obtain the posterior draw of $(m^{s}_{\eta}(X_{i}))_{i=1}^{n}$ using the Gaussian process prior and the data from the control arm.
@@ 3.2 A Double Robust Bayesian Method
Our Bayesian approach relies on prior correction via inverse propensity score weighting in the least favorable direction as specified by the efficient influence function. In contrast to Abadie (2005), we do not incorporate inverse propensity score weighting (IPW) directly. Importantly, we make use of IPW for the prior correction of our Bayesian procedure. This resembles Sant’Anna and Zhao (2020) who combine OR and IPW to achieve doubly robust estimation in the frequentist setting. Following Hahn (1998); Hirano, Imbens, and Ridder (2003) or, in the DiD setup Sant’Anna and Zhao (2020), the efficient influence function for the ATT is given by
@@
We show in the Supplemental Appendix B that the Riesz representer $\gamma_{\eta}$ determines the least favorable direction associated with the Bayesian submodel with the largest variance. Our prior adjustment using this Riesz representer provides exact invariance under shifts in nonparametric components along this direction. This extends the work of Ray and van der Vaart (2020) on unconditional average treatment effects to the ATT case, where the least favorable function, such as the efficient influence function, takes a different functional form. In a similar vein to Breunig, Liu, and Yu (2023), we use the Riesz representer to correct for posterior bias under double robust smoothness conditions.
@@
     (a) Obtain the posterior draw of $(m^{s}_{\eta}(X_{i}))_{i=1}^{n}$ using the adjusted prior in (3.7) and the data from the control arm
@@
Also the double robust Bayesian Algorithm 2 allows for simultaneous point estimation and uncertainty quantification. Our $100\cdot(1-\alpha)\%$ credible set $\mathcal{C}_{n}(\alpha)$ for the ATT parameter $\tau_{0}$ is as in (3.3), but here $q(a)$ denotes the $a$ quantile of $\{\check{\tau}_{\eta}^{s}:s=1,\ldots,B\}$. The Bayesian point estimator by $\overline{\tau}_{\eta}=B^{-1}\sum_{s=1}^{B}\check{\tau}_{\eta}^{s}$. For implementation of the prior process $W^{m}$, we make use of a zero-mean Gaussian process with the commonly used squared exponential (SE) covariance function (Rassmusen and Williams, 2006, p.83). Specifically, the covariance function is determined by $K\left(x,x^{\prime}\right):=\nu^{2}\exp\left(-\sum_{l=1}^{p}a_{ln}^{2}(x_{l}-x^{\prime}_{l})^{2}/2\right)$, where the hyperparameter $\nu^{2}$ is the kernel variance and $a_{1n},\ldots,a_{pn}$ are rescaling parameters that reflect the relevance of treatment and each covariate in predicting $\eta^{m}$. In practice, they can be decided by maximizing the marginal likelihood. Regarding the pilot estimator for the propensity score, we employ logistic regression. As a pilot estimator of the conditional mean function, we use the posterior mean of standard Gaussian process priors following Breunig, Liu, and Yu (2023). To be specific, $\widehat{m}(x)=\sum_{s=1}^{B}m_{\eta}^{s}(x)/B$, where $m_{\eta}^{s}$ is obtained in Step (a) of posterior computation in Algorithm 1.
@@
The correction term in (3.9) corrects the posterior distribution by an adjustment term which is required under double robustness as we show in Theorem 5.1. Using this posterior correction, Theorem 5.2 establishes BvM results under double robust smoothness assumptions. On the other hand, posterior correction is not required if one is willing to impose Donsker type smoothness conditions on the conditional mean function $m_{\eta}$, i.e., if the smoothness of $m_{\eta}$ exceeds $\dim(X_{i})/2$, which is an implication of Corollary 5.1.
Posterior correction was also proposed by Breunig, Liu, and Yu (2023) in the context of average treatment effects (ATEs) using cross-sectional data. In their case, the bias correction term is given by $\widehat{b}^{ATE,s}_{\eta}=n^{-1}\sum_{i=1}^{n}\bm{\tau}\left[m_{\eta}^{s}-\widehat{m}\right]\left(Z_{i}\right)$, where $\bm{\tau}[m](z):=m(1,x)-m(0,x)+\widehat{\gamma}^{ATE}(d,x)(y-m(d,x))$ is an estimator of the efficient influence function for the ATE. See also Remark 5.1 for the a more explicit comparison of the biases in both cases. We observe that double-robust Bayesian inference for the ATT, as given in (3.9), involves a simpler form of posterior correction compared to that for the ATE.

### v2

@@ Abstract
Keywords: Difference-in-differences, conditional parallel trends, semiparametric Bayesian inference, Bernstein–von Mises theorem, double robustness, semiparametric efficiency.
@@ 1 Introduction
We also provide an extension of our Bayesian procedure, which incorporates robustification via estimated propensity scores and is particularly suited for more complex models, either due to a larger number of continuous covariates or when the underlying conditional mean functions are not smooth. Our Double Robust Bayesian procedure adjusts the prior and posterior distributions by incorporating an efficient influence function. By doing so, we leverage the rich frequentist literature on double-robust estimation, specifically Sant’Anna and Zhao (2020) in the DiD framework, without sacrificing many of the desirable properties of the Bayesian approach. Under double-robust smoothness conditions, our robust Bayesian procedure satisfies the semiparametric Bernstein–von Mises (BvM) theorem, albeit with a “bias term” in the posterior. Specifically, the resulting posterior distribution depends on the unknown true conditional mean and propensity score functions. Our double-robust Bayesian approach addresses this “bias term” by incorporating an explicit posterior correction. Both the prior adjustment and the posterior correction are derived from functional forms closely associated with the efficient influence function of estimating the ATT. Due to differences in the efficient influence functions for average treatment effects (ATE) and ATT, our double-robust method shares similarities with but differs in functional form corrections from Breunig, Liu, and Yu (2023)’s work on ATE. Notably, we find that the bias term in the BvM for ATT is substantially simpler than that for ATE, which also explains the favorable finite-sample behavior of Bayesian ATT estimators even without posterior corrections.
@@
In our Monte Carlo simulations, we find that our methods result in improved empirical coverage probabilities while maintaining competitive confidence interval lengths compared to existing frequentist methods. This finite sample advantage is also observed in low dimensional cases for our Bayesian method that does not involve prior or posterior corrections. [^1] In contrast, a Bayesian method without prior correction performs poorly for the average treatment effect (ATE) as shown by Breunig, Liu, and Yu (2023). This can be explained by the construction of the Bayesian procedure for the ATT, which involves only a prior specification for the conditional mean function in the control arm. In particular, we note that our approach leads to more accurate uncertainty quantification and is less sensitive to estimated propensity scores that are close to boundary values. Our Bayesian methodology requires a prior specification through a likelihood function for the control arm, for which we impose an exponential family structure. In the Gaussian case, for instance, this leads to a computationally efficient procedure with the posterior being multivariate Gaussian, which avoids computational demanding methods like MCMC. We stress that the misspecification of this structure does not have serious consequences for estimation of the ATT. First, as shown in (Kleijn and van der Vaart, 2006, Section 4), nonparametric Bayesian methods possess the same robustness to misspecification as the frequentist M-estimation using least squares. Second, we provide finite sample evidence through simulations, where we find that our Bayesian procedures are not sensitive to misspecifications of the likelihood functions.
@@
The remainder of this paper is organized as follows. Section 2 presents the setup and introduces the Bayesian framework in the DiD setup. Section 3 outlines our Bayesian methods. In Section 4, we establish inference via semiparametric BvM theorems for our first method. In Section 5, we derive a doubly robust, semiparametric BvM theorem for our second method. Section 6 provides BvM results under primitive conditions when using squared expontential process priors. Section 7 presents finite sample results via simulations and an empirical illustration. Towards the end, we outline two extensions of our methodology: Section 8 provides an extension to the staggered intervention with multiple time periods and repeated cross-sectional data. Proofs of main theoretical results are collected in Appendix A. Supplementary Appendices BE provide additional technical results and further simulation evidence.
@@ 2.1 Setup
the conditional mean of the differenced outcome across two periods, where $\Delta Y_{i}:=Y_{i2}-Y_{i1}$. The researcher observes an independent and identically distributed (i.i.d.) observations of $(Y_{i1},Y_{i2},D_{i},X_{i}^{\top})^{\top}$, $i=1,\dots,n$. In addition to this canonical panel data setup, we discuss how our results translate to repeated cross-sections data in Section 8.2, while an extension to staggered intervention is provided in Section 8.1. In addition to this canonical panel data setup, we provide an extension to staggered interventions in Section 8.1 and discuss how our results translate to repeated cross-sectional data in Section 8.2. For notational simplicity, we will henceforth suppress the unit index $i$.
@@
One can construct an estimator that replaces the conditional mean function $m_{0}$ with an estimator, known as the outcome regression approach, as described in Heckman, Ichimura, and Todd (1997). As noted by (Abadie, 2005, p.6), plug-in estimators based on standard nonparametric estimators of the conditional mean function $m_{0}$ can face significant challenges due to the curse of dimensionality. This is where we can capitalize the strength of Bayesian estimation, which allows us to incorporate rich covariate information in the prior distribution. It has also been noted in the recent literature that, in the presence of heterogeneous treatment effects in $X$, i.e., when $\mathbb{E}_{0}[Y_{2}(1)-Y_{2}(0)\mid X=x,D=1]$ varies with $x$, the two-way fixed effect estimator (TWFE) is in general not consistent for the ATT. See also Remark 1 of Sant’Anna and Zhao (2020) for an explicit discussion.
@@ 2.2 A Bayesian Framework
where $f_{\Delta Y\mid D,X}(y\mid d,x)$ for $d=1$ is the unrestricted conditional density of $\Delta Y$ given $(D,X)$, while for $d=0$, we impose the exponential family condition as in (2.4). Here, $f$ denotes the joint density of $(D\Delta Y,D,X^{\top})^{\top}$ under $P_{\eta}$, and the corresponding cumulative distribution function is denoted by $F$. Importantly, specifying only the prior distribution on density function $f$ and the conditional mean function $m$ is sufficient for identifying the ATT parameter of interest. Specifically, there is no need to additionally parameterize the propensity score $\pi$, the marginal density of $X$, or the conditional density for the control group, $f_{\Delta Y|D,X}(y\mid 0,x)$, which is a deterministic function of $m(x)$ due to the exponential family assumption imposed in (2.4) below.
We consider the following reparametrization of $(m,f)$ given by $\eta=(\eta^{m},\eta^{f})$. A central insight of this paper is to show that a nonparametric process prior specification on the conditional mean function $m$ and the density $f$ is sufficient for the Bayesian inference on the ATT parameter. We index the probability model by $P_{\eta}$, where
$m_{\eta}=q^{-1}(\eta^{m})\,\,\text{ and }\,\,f_{\eta}=\exp(\eta^{f})$
for some known, invertible function $q(\cdot)$, which we specify below. We can write the ATT depending on a hyperparameter $\eta$ as
@@
The family (2.4) allows for counting and continuous outcomes. For instance, when $a=1$, the Poisson distribution corresponds to the choices $c(y)=1/(y!)$, $q(m)=\log m$, and $A(m)=m$, while the exponential distribution is represented by $c(y)=1$, $q(m)=-1/m$, and $A(m)=\log m$. Furthermore, the normal distribution with $\text{Var}(\Delta Y|D=0,X)=\sigma^{2}$ for some $\sigma>0$, is captured by $c(y)=\exp(-y^{2}/(2\sigma^{2}))/\sqrt{2\pi\sigma^{2}}$, $q(m)=m/\sigma$, $A(m)=m^{2}/(2\sigma^{2})$, and $a=1/\sigma$. For the normal case, we treat $\sigma$ as a hyperparameter and estimate it together with hyperparameters in Gaussian process prior by maximizing marginal likelihood. We note that while a generalization to multinomial outcomes, as in Breunig, Liu, and Yu (2023), is possible, we do not consider this case explicitly in this paper.
@@ 3.1 Bayesian Inference using Standard Gaussian Process Priors
We first consider nonparametric Bayesian inference, which builds on a standard Gaussian process prior for the conditional mean function combined with an independent Dirichlet process prior for the conditional expectation in (2.3). The proposed method does not include a propensity score adjustment, which prevents it from achieving double robustness, as we see in the next section. However, simulations show that the proposed method is robust to a failure of overlap.
@@
The Dirichlet process is default prior on spaces of probability measures. By the definition of the ATT $\tau_{\eta}$, we assign a Dirichlet process prior to model the distribution $F_{\eta}$, which induces the so-called Bayesian bootstrap when the base measure of the Dirichlet process is taken to be zero; see (Rubin, 1981) and also Chamberlain and Imbens (2003).
@@
     (a) Generate the $s$-th draw of the posterior of $(m_{\eta}(X_{i}))_{i=1}^{n}$ using the Gaussian process prior and the data from the control arm; denote it as $(m^{s}_{\eta}(X_{i}))_{i=1}^{n}$.
@@ 3.2 A Double Robust Bayesian Method
Our Bayesian approach relies on prior correction via inverse propensity score weighting (IPW) in the least favorable direction, as specified by the efficient influence function. In contrast to Abadie (2005), we do not incorporate IPW directly. Importantly, we make use of IPW for the prior and posterior correction of our Bayesian procedure. This resembles Sant’Anna and Zhao (2020) who combine OR and IPW to achieve doubly robust estimation in the frequentist setting. Following Hahn (1998); Hirano, Imbens, and Ridder (2003) or, in the DiD setup Sant’Anna and Zhao (2020), the efficient influence function for the ATT is given by
@@
We show in the Supplemental Appendix B that the Riesz representer $\gamma_{\eta}$ determines the least favorable direction associated with the Bayesian submodel with the largest variance. Our prior adjustment using this Riesz representer provides exact invariance under shifts in nonparametric components along this direction. This extends the work of Ray and van der Vaart (2020) on unconditional average treatment effects to the ATT case, where the least favorable function, given by the efficient influence function, takes a different functional form. In a similar vein to Breunig, Liu, and Yu (2023), we use the Riesz representer to correct for posterior bias under double robust smoothness conditions.
@@
     (a) Generate the $s$-th draw of the posterior of $(m_{\eta}(X_{i}))_{i=1}^{n}$ using the adjusted prior in (3.7) and the data from the control arm; denote it as $(m^{s}_{\eta}(X_{i}))_{i=1}^{n}$.
@@
Also the double robust Bayesian Algorithm 2 allows for simultaneous point estimation and uncertainty quantification. Our $100\cdot(1-\alpha)\%$ credible set $\mathcal{C}_{n}(\alpha)$ for the ATT parameter $\tau_{0}$ is as in (3.3), but here $q(a)$ denotes the $a$ quantile of $\{\check{\tau}_{\eta}^{s}:s=1,\ldots,B\}$. The Bayesian point estimator by $\overline{\tau}_{\eta}=B^{-1}\sum_{s=1}^{B}\check{\tau}_{\eta}^{s}$. For implementation of the prior process $W^{m}$, we make use of a zero-mean Gaussian process with the commonly used squared exponential (SE) covariance function (Rassmusen and Williams, 2006, p.83). Specifically, the covariance function is determined by $K\left(x,x^{\prime}\right):=\nu^{2}\exp\left(-\sum_{l=1}^{p}a_{ln}^{2}(x_{l}-x^{\prime}_{l})^{2}/2\right)$, where the hyperparameter $\nu^{2}$ is the kernel variance and $a_{1n},\ldots,a_{pn}$ are rescaling parameters that reflect the relevance of treatment and each covariate in predicting $\eta^{m}$. In practice, they can be chosen by maximizing the marginal likelihood. Regarding the pilot estimator for the propensity score, we employ logistic regression. As a pilot estimator of the conditional mean function, we use the posterior mean of standard Gaussian process priors following Breunig, Liu, and Yu (2023). To be specific, $\widehat{m}(x)=\sum_{s=1}^{B}m_{\eta}^{s}(x)/B$, where $m_{\eta}^{s}$ is obtained in Step (a) of posterior computation in Algorithm 1.
@@
A posterior debiasing step for the posterior is required by Theorem 5.1 and, as shown in Theorem 5.2, our posterior correction term in (3.9) indeed allows for a derivation of the BvM result under double robust smoothness assumptions. On the other hand, the posterior correction is not required if one is willing to impose Donsker type smoothness conditions on the conditional mean function $m_{\eta}$, i.e., if the smoothness of $m_{\eta}$ exceeds $\dim(X_{i})/2$, which is an implication of Corollary 5.1.
Posterior corrections were also proposed by Breunig, Liu, and Yu (2023) in the context of average treatment effects (ATEs) using cross-sectional data. In their case, the bias correction term is given by $\widehat{b}^{ATE,s}_{\eta}=n^{-1}\sum_{i=1}^{n}\bm{\tau}\left[m_{\eta}^{s}-\widehat{m}\right]\left(Z_{i}\right)$, where $\bm{\tau}[m](z):=m(1,x)-m(0,x)+\widehat{\gamma}^{ATE}(d,x)(y-m(d,x))$ is an estimator of the efficient influence function for the ATE. See also Remark 5.1 for the a more explicit comparison of the biases in both cases. We observe that double-robust Bayesian inference for the ATT, as given in (3.9), involves a simpler form of posterior correction compared to that for the ATE.